INFRARED SPECTROSCOPY OF JET-COOLED TROPOLONE AND TROPOLONE-OD IN THE O-H AND C-H STRETCH REGION

\underline{C. A. ARRINGTON}, R. K. FROST, F. C. HAGEMEISTER, AND T. S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Fluorescence-dip infrared spectroscopy has been employed to record the vibrational spectra of gas phase tropolone and tropolone-OD in the O-H and C-H stretch region for the first time. Tropolone, an pseudoaromatic ringed system, possesses a symmetric O^{..}H − O tunneling coordinate which produces a ground state tunneling splitting of 0.99 cm^{−1}. Use of a two laser depletion scheme allows separate infrared spectra to be recorded for both the lower (a_{1} symmetry) and upper (b_{2} symmetry) ground state tunneling levels. In each spectrum, a single transition ascribable to the O-H stretch is observed. The (a_{1}) O-H stretch at 3134.9 cm^{−1} is close in energy to the (b_{2}) O-H stretch observed at 3133.9 cm^{−1} indicating that the transitions are to the same excited state tunneling level. Tropolone-OD was investigated in order to identify the C-H stretch vibrations which are close in energy to the red-shifted O-H stretch of tropolone. The spectroscopic results along with recent ab initio calculations are discussed in terms of the hydrogen tunneling motion in tropolone.